To produce 80 copies of a magazine costs 52 cents per copy.
To produce 760 copies of the magazine costs 35 cents per copy.
a) Name two ordered pairs.
b) Write the linear equation.
c) Predict the cost per copy for 480 copies.

Answer by AnlytcPhil(1116)

To produce 80 copies of a magazine costs 52 cents per copy. To produce 760
copies of the magazine costs 35 cents per copy.
a) Name two ordered pairs.
(80,52) and (760,35)
b) Write the linear equation.

`
Use the slope formula:
` ` `y2 - y1
m = ---------
` ` `x2 - x1
`
with (x1,y1) = (80,52) and (x2,y2) = (760,35)
`
` ` ` 35 - 52 -17 -1
m = --------- = --- = --- = -0.025
` ` `760 - 80 680 40
`
Now use the point-slope formula:
y - y1 = m(x - x1)
y - 52 = -0.025(x - 80)
y - 52 = -0.025x + 2
` ` `y = -0.025x + 54

`
c) Predict the cost per copy for 480 copies.
Substitute 480 for x in that equation
` ` `y = -0.025(480) + 54
` ` `y = 42
`
Edwin J